Find all solutions:
3+2 log(x+5)=5

Answer by edjones(8007)

3+2*log(x+5)=5
2*log(x+5)=2
log(x+5)^2=2
10^log(x+5)^2=10^2
(x+5)^2=100
x^2+10x+25-100=0
x^2+10x-75=0
(x+15)(x-5)=0
x=5
.
Ed